OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned? 44 40 76 80

OpenStudy (texaschic101):

w + m = 120 --> w = 120 - m 2m - 3w = 20 now sub 120 - m in for w in the second equation

OpenStudy (anonymous):

lets say men are x and women are y then twice the number of men wil become 2x and three times the number of women becomes 3y right? so according to the given condition: "twice the number of men less three times the number of women is equal to 20" it will become 2x-3y = 20. ________________ equation 1 and we have x + y = 120 ._________________equation 2. solving you get x and y, he asked for x.